OpenStudy (anonymous):

if sin(x)=1/3, what is/how do you find sin(-x)?

jimthompson5910 (jim_thompson5910):

hint: sine is an odd function which means sin(-x) = -sin(x)

jimthompson5910 (jim_thompson5910):

all odd functions have the property that f(-x) = -f(x)

OpenStudy (phi):

one way is with a picture, and the definition of sin(x) = opp/hypotenuse  -x means go *clockwise* from the positive x-axis by symmetry we see the opposite side is -1, and the hypotenuse is still +3 and sin(-x)= -1/3